Rafael must choose a number between 67 and 113 that is a multiple of 3, 6, and 7. Write all the numbers that he could choose. If there is more than one number, separate them with commas.
Answer by ikleyn(52830)

.
The multiples of 3, 6 and 7 are 6*7= 42, 2*42 = 84, 3*42 = 126 and so on.

You do the rest . . .